Fermi National Laboratory
 

Accelerator Update for Jun 6, 2006 - June 7, 2006

 

Tuesday June 6

The day shift began with the Main Injector (MI), Antiproton Source (Pbar), Recycler, Switchyard (SY), NuMI, MiniBooNE, and the Tevatron (TeV) conducting studies, and with NuMI taking beam.

TeV experts loaded a 12x0 study store at 8:46 AM.

Operations had the Collider Detector at Fermilab (CDF) and D-Zero (D0) experimental halls ready for access by 9:41 AM.

Shot setup for transfers of antiprotons from Pbar to the Recycler began at 9:57 AM.

The first few transfers of antiprotons didn’t make it into the Recycler, but the fourth through the eleventh transfers did. Pbar resumed stacking at 12:45 PM.

Experts began commissioning the Pbar/NuMI mixed mode stacking at 1:52 PM.

Pbar experts began a Debuncher phasing study at 2:29 PM.

MI went back to separate events for Pbar and NuMI beam at 9:12 PM.

At 9:12 PM, the Klystron Debuncher tripped on a cavity LCW flow indication for the first of five times in three and a half hours. It reset each time.

At 11:45 PM, Operations reported problems running mixed mode. Operators consulted with an expert and then checked the dampers. Later they discovered that they had lost the beam permit due to a water skid that wouldn’t reset.

Wednesday June 7

The midnight shift began with Operations monitoring the stash, with Pbar stacking, with the TeV conducting chromaticity studies, with MI, RR, and Switchyard (SY) conducting studies, and with MiniBooNE and NuMI taking beam.

At 12:28 AM, the Klystron Debuncher tripped on a LCW flow indication again. It reset okay.

At 2:05 AM, Operations reported that they called an expert about the 400 MeV Klystron Debuncher water skid trips. The expert instructed operators to increase the system’s pressure.

At 3:54 AM AM, the Klystron Debuncher tripped on a LCW flow indication again. It reset okay.

Operations began, at 6:47 AM, preparing the D0 and CDF halls for access.

Operations prepared the Recycler at 7 AM to stash. The stash ended at 7:25 AM.

Operations prepared the TeV for access at 7:16 AM.

Operations prepared Pbar for access at 7:39 AM.

The Plans for Wednesday

The plan is to allow D0 its sixteen-hour access. CDF will make a shorter access. There will be a transfer hall access to install the A0 quadrupole magnet for SY. Several other machines will make short accesses.

Machine Reports

Linac

Experts fixed a 400 MeV low-level RF problem.

Booster

Experts conducted orbit work and tuned the high order correctors.

Main Injector

Experts worked on mixed mode slip stacking.

Antiproton Source

Experts reported that the first transfer of antiprotons was good, but there are problems that need fixing. They need to tune up stacking; the stacking rate has fallen off since Monday. Experts also want to make some lattice measurements on Thursday or Friday.

TeV

Experts continue to commission the machine. Their tune of the central orbit is on schedule. Experts also worked on coupling and chromaticity measurements.

Recycler

Experts have seen antiprotons and electrons in their machine. They need to work on energy matching. The experts are ready to send antiprotons back to MI.

Switchyard 120

Experts tested power supplies. The A0 quadrupole magnet will be installed today during the D0 access. They would like to have extracted beam on Thursday.

MiniBooNE

The experiment is taking data.

NuMI

Experts reported that the mixed mode at 2 seconds was working. They also reported that the Horn cooling system was fixed.

Collider Detector at Fermilab

They will access their hall today and work on an intermittent readout problem.

D-Zero

Experts fixed their collimator and many other minor problems yesterday. They will access their hall today to make repairs.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/07/2006    email Fermilab